As part of your work on the audit of accounts receivable, you are also tasked with evaluating the reasonableness of Black Panther’s allowance for doubtful accounts of $15,500. Your audit firm has set performance materiality at $20,000. To assist you with your evaluation, the client has provided an aging of accounts receivable schedule and some details about a few specific customer accounts. Using this information, provide your best estimate for the allowance for doubtful accounts and propose an adjusting entry if you think one is needed.1 0 -30 $565,836 1 Days Old 31 - 60 | 60 - 90 $181,100 $43.252 90 – 120 $3.555 > 120 $1.500 Balance

Company policy requires that 10% of accounts between 31 – 60 days have a reserved established, 20% of accounts between 60 – 90 days, 60% of accounts 90 – 120 days, and 95% of accounts over 120 days.

In addition, the client noted that one customer, Grande Enterprises experienced a catastrophic loss and is experiencing cash flow problems. Grande has assured Black Panther they will pay their bills in full. Grande’s entire account balance of $21,000 is between 31 – 60 days outstanding. Additionally, Minaj Holdings has disputed two invoices from its customer account. The first is for $3,555 and is 92 days outstanding. The second is for $8,000 and is only 15 days outstanding. Minaj states that they will not make any payments on either invoice ever. Black Panther believes its in-house collection team will get at payment for both invoices despite its collection department historically only collecting about 45% of outstanding bills.

I AM NOT JUST LOOKING FOR THE ANSWER, I WOULD LIKE SOMEBODY TO PLEASE EXPLAIN THE PROCESS OF HOW THE ANSWER IS DERIVED AS WELL AS ALL NUMBERS AND COMPUTATIONS PLEASE
Answer #1

Black Panthor's Reserve for doubtful debts Policy - Black Panthor's Policy of recognising allowance for doubtfull debts in relation to receivable's ageing looks correct as the percentage increases as the receivable becomes more older for instance reserve is created at 95% for the accounts outstanding since 120 days.

Calculation of allowance for doubtfull accounts - To Proceed with the calculation of allowance we will need to first bifercate total receivables into two categories - A. First Category - Debtors where no disputs are pending B. Second Category - Debtors with payment disputes. We will sum up outcomes of A and B to arrive at the final allowance amount.

A. First Category

Days Balance (a) Disputed amount (b) Normal Receivables (a-b) Reserve
0-30 565836 8000 557836 0 (0%)
30-60 181100 0 181100 18110 (10%)
60-90 43252 0 43252 8650.40 (20%)
90-120 3555 3555 3555 2133 (60%)
>120 1500 0 1500 1425 (95%)

Total of the column Reserve is 30318.4.

B. Second Category

Though Grande Enterprises experienced a catastrophic loss, the receivables are older of 30-60 days and Grade has also assured full payment. As an Auditor checking of communication in writing with Grande shlould form part of audit checklist while performimg receivables audit. If audit evidence shows a reliance can be placed on Grande's statement, No reserve to be created.

In case of Minaj Holdings, In house team will recover upto 45% and hence on the basis of conservatism full reserve to be created for the receivables irrespective of days of outstanding. Auditor should consider all the facts and information which he come accross in audit and audit materiality can be reset if required.  Reserve for Minaj = 11555 * 65% = 7510.75

Total Reserve to be created will be 37,829.15 ( A + B = 30318.4 + 7510.75 ). Black Panther’s allowance for doubtful accounts of $15,500 is not in line with the policy and additional information. Hence Additional allowance of Rs. 22,329.15 need to be created.